let f(x) be a 4 th degree polynomial such that f(1)= 1 , f(2)=1,f(3)=1,f(4)=1 , find f(5)

Grade:11

1 Answers

Arun
6 years ago
All has value equal to 1.
Hence we try to make a polynomial which has value 1
when x = 1, 2, 3, 4
Hence
Let f(x) = (x-1)(x-2)(x-3)(x-4) +1
Hence
f(5) = 4*3*2*1 +1 = 25